Melanie feeds her cat gourmet cat food that cost $1.75 per pound. She combines it with cheaper food that cost $0.50 per pound. How many pounds of cheaper food should Melanie buy to go with 5 pounds of gourmet food, if she wants the average price to be $1.00 per pound?
Answer by amoresroy(361)

Let x = number of pounds of cheaper food Melanie shoud buy...
Equation:
1.75 * 5 + 0.5x = 1.00*(5+X)
8.75 + 0.5x = 5 + x
Combine like terms
0.5x = 3.75
Muliply by 2 to get x
x = 7.5
Melanie should buy 7.5 pounds of cheaper food